OpenStudy (anonymous):

solve each system by substiution y=4x-9 y=x-3

OpenStudy (anonymous):

4x-9=x-3 3x=6 x=2

OpenStudy (anonymous):

4x-9=x-3 3x=6 x=2

OpenStudy (anonymous):

y=2-3 y=-1

OpenStudy (anonymous):

\[\color{salmon}{\Huge{\checkmark}}\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

lol I just wanted to try it. : )

hero (hero):

Ishaan, you stole my checkmark! lol

OpenStudy (anonymous):

lol I did !